• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ole Hansen

    Log entry time 15:51:24 on April10,2007

    Entry number 198490

    keyword=Day shift summary

    Main events
    08:00 Start shift ready for data taking at Q2 = 55 kinematics. Waiting for beam, which is being tuned for Hall B.
    08:45-09:45 Harp scans. Eventually determine beam size to be about 65x175 uA, which is very good.
    09:50-11:50 Production running on 3He cryo cell. Quick replay looks decent, although statistics are low in left arm and for coincidences.
    11:50-16:00 Beam gone for Hall B tuning.

    Waiting for beam with ETA later tonight at end of shift.

    Run summary
    2281 - test run during harp scans, probably junk.
    2282-2287 3He production, DT=9%, 70uA, 1M events each.